OpenStudy (anonymous):

Solve the given equation,rational equations. 29=(3/s)+12

OpenStudy (anonymous):

\[29 = \frac{ 3 }{ s } + 12\] Bring 12 to the other side. \[29 - 12 = \frac{ 3 }{ s } => 17 = \frac{ 3 }{ s }\]

OpenStudy (anonymous):